OpenStudy (anonymous):

Mars has an approximate diameter of 6.794 x 103 kilometers. Assume that Mars is a perfect sphere. Use the formula for the volume of a sphere, V = 4/3 πr3 to determine the volume of Mars. Show your work and give the answer in scientific notation.

OpenStudy (helder_edwin):

d=2r

OpenStudy (anonymous):

solve for r and then plug r into the formula for the volume of a sphere.

OpenStudy (ash2326):

@kenneyfamily did you understand?

OpenStudy (anonymous):

no sorry..still confused

OpenStudy (ash2326):

@Romero would you please help here?

OpenStudy (anonymous):

well we know that the diameter can also be rewritten as d=r2 where r=radius and d=diameter and we know that the only unknown in the formula is r

OpenStudy (anonymous):

Now let's solve for r 2r=d r=d/2

OpenStudy (anonymous):

so all we have to do is divide the diameter by 2 (6.794 x 103 kilometers)/2=r

OpenStudy (anonymous):

\[\frac{4}{3} πr^3\] where r=(6.794 x 103)/2

OpenStudy (anonymous):

just plug that in into a calculator and you get the answer for volume

OpenStudy (ash2326):

@Romero good explanation :)

OpenStudy (anonymous):

THANKS!!